OpenStudy (anonymous):

what is the solution to the following 3x + 3y = 10 -9x - 9y = -30

OpenStudy (kc_kennylau):

-3*(1)=(2) therefore infinitely many solutions

OpenStudy (anonymous):

3(3x+3y = 10) gives 9x + 9y = 30 -9x-9y=-30 we're eliminating here 0 = 0 You have an infinite number of solutions that will work.

OpenStudy (wolf1728):

The second equation is just the first equation multiplied by -3, hence they are the same equation and you'll have an infinite number of solutions.

OpenStudy (sleepyjess):

3x + 3y = 10 3x = -3y+10 \[x=\frac{ -3y+10 }{ 3 }\]Then substitute this into the 2nd equation